# Three-dimensional gravity-capillary standing waves: computation, resonance and instability

### v1

@@ 5 Temporal evolutions
$\widehat{p}=\widehat{\eta}+\frac{\mathrm{i}\widehat{\varphi}}{\sqrt{|\boldsymbol{k}|+1/|\boldsymbol{k}|}},\quad\widehat{q}=\widehat{\eta}-\frac{\mathrm{i}\widehat{\varphi}}{\sqrt{|\boldsymbol{k}|+1/|\boldsymbol{k}|}},$ (57)
and rewrite (55) and (56) as
@@
Figure 31: Temporal evolution of a non-resonant Case II standing wave $(k=l=1,H=0.29)$ perturbed by $0.001\cos(x+y)$. (a) Time histories of $|\widehat{p}_{1,1}|$ and $|\widehat{p}_{-1,1}|$. The black dashed lines label the instants when the curves reach maximums and minimums. (b)-(f) Representative surface profiles at $t=0,72T,137.1T,204.1T$ and $273.2T$.
@@
Figure 32: Temporal evolution of a non-resonant Case II standing wave $(k=l=1,H=0.5)$ perturbed by $0.001\cos(x+y)$. (a) Time histories of $|\widehat{p}_{1,1}|$ and $|\widehat{p}_{1,-1}|$. The black dashed lines label the instants when the curves intersect and reach maximums and minimums. (b)-(c) Representative surface profiles at $t=26T$ and $61.6T$.
@@
Next we examine the stability of three-dimensional standing waves. Zhu et al. [zhu2003three] have demonstrated that two-dimensional gravity standing waves, say, in $x$-direction with a wavenumber $\kappa$, are linearly unstable to harmonic disturbances $\mathrm{e}^{\mathrm{i}(\alpha x+\beta y)}$ if $\alpha^{2}+\beta^{2}\approx\kappa^{2}$. From a point of view of resonance, this condition implies that the disturbances have approximately the same frequency as the base standing waves, thereby supporting energy transfer between them. Over long periods, they showed that the instability leads to nearly cyclic return to the initial condition. Suppose a similar instability mechanism exists in gravity-capillary case, then (1) implies that the three-dimensional standing waves are likely unstable as well. We start to examine this idea by considering a non-resonant Case II standing wave with $k=l=1,H=0.29$, as shown in 31 (b). To satisfy the periodic boundary condition, we initially perturb this solution by a small disturbance $0.001\cos(x+y)$, which corresponds to two complex harmonics $0.0005\mathrm{e}^{\pm\mathrm{i}(x+y)}$. Figure 31 (a) shows the time histories of $|\widehat{p}_{1,1}|$ and $|\widehat{p}_{1,-1}|$ at integer temporal periods, which can be regarded as ensemble averages measuring the typical amplitudes of $\cos(x+y)$ and $\cos(x-y)$ modes. These two curves clearly demonstrates that energy is transferred between the two modes in a quasi-periodic fashion with an approximate period of $70T$. Therefore, the base standing wave is unstable to the initial perturbation. Panels (c)-(f) show four representative surface profiles at $t=71T,137.1T,204.1T$ and $273.2T$. Together with the initial profile in (b), these form a near cyclic return to the initial standing wave. For base standing waves with larger amplitude and the same initial disturbance, we observe similar cyclic returns occurring on relatively shorter temporal periods. Furthermore, there could be two alternatively appearing quasi-two-dimensional modes. A temporal evolution initiated with a Case II standing waves for $k=l=1,H=0.5$ is shown in figure 32. Panel (a) presents the curves of $|\widehat{p}_{1,1}|$ and $|\widehat{p}_{1,-1}|$ taken at each integer period, and panels (b) and (c) correspond to the surface profiles at $t=26T$ and $61.6T$ when the quasi-two-dimensional modes appear.
@@
Figure 33: Temporal evolution of a branch-$1$ standing wave $(k=1,l=1.5312,H=0.05)$ perturbed by $0.001\cos(x+1.5312y)$. (a) Time histories of $|\widehat{p}_{1,1.5312}|$, $|\widehat{p}_{1,-1.5312}|$, and $|\widehat{p}_{0,3.0624}|$. (b)-(c) Representative surface profiles at $t=0$ and $453.2T$.
@@
In general, consider three-dimensional standing waves dominated by $\cos(kx)\cos(ly)\cos(\omega t)+\cos(lx)\cos(ky)\cos(\omega t)$ and perturbed by $\delta\cos(kx\pm ly),\delta\ll 1$. It is found that such an instability is universal and the resulting energy-transfer rate is related to the frequencies of the base standing wave and the one used as initial perturbation. For two dimensional standing waves with wavenumber $\kappa=\sqrt{k^{2}+l^{2}}$, Concus [concus1962standing] showed that the frequency $\omega$ is a monotonically decreasing function of wave amplitude unless $\kappa\in(\kappa_{1},\kappa_{2})$. For infinite deep water, $\kappa_{1}=0.5154$ and $\kappa_{2}=\sqrt{2}/2$ (see figure 1 in [concus1962standing]). Therefore, if $\kappa$ lie outside this specific interval and the frequency curves of the base standing waves have down-shifts close to the bifurcation point, the base standing waves are likely unstable. Otherwise, there is no overlap between the the frequency curves of the base stadning waves and the perturbations, admitting no significant energy transfer. To confirm this, we consider the branch-$1$ standing waves with $k=1,l=1.5312$ whose the leading Fourier components can be decomposed into
@@
We first use the initial disturbance $0.001\cos(x+1.5312y)$ whose frequency is close to $2.8188$. As shown in figure 23, the frequency curve of the base standing wave is slightly above this value for $H<0.12$. Figure 33 exhibits the temporal evolution of the standing wave with $H=0.05$ whose frequency is $\omega=2.8224$. Panel (a) shows the time histories of $|\widehat{p}_{1,1.5312}|,|\widehat{p}_{1,-1.5312}|$, and $|\widehat{p}_{0,3.0624}|$, which are taken at integer temporal periods and represent the typical amplitudes of $\cos(x+1.5312y)$, $\cos(x-1.5312y)$, and $\cos(3.0624y)$ modes. Energy transfer among the three modes is weak and only has influences over long periods. Note that the evolution also demonstrates a nearly periodic return to the initial state (with a phase shift) with an approximate period of $550T$, as shown by the surface profiles in panels (b) and (c).
@@ 6 Conclusions
We showed that the cubic and quintic models align closely with the full-potential formulation in both two and three dimensions. Typical non-resonant gravity-capillary standing waves exhibit round crests and troughs, in contrast to the pyramidal shape of pure gravity standing waves. When resonances occur, standing waves exhibit complicated and rapidly oscillating surface patterns, and their bifurcations could break into multiple, disjoint solution branches. In particular, we considered the resonant standing waves that support the three-wave resonance, generalising the concept of classical Wilton ripples occurring in two-dimensional gravity-capillary travelling waves. For both collinear and non-collinear resonant triads, we identified the existence of multiple solution branches bifurcating from the solutions, along with various forms of resonant standing waves featuring square, hexagonal, and more complex flower-like surface patterns. We examined the temporal periodicity of the computed standing waves via initial-value calculations, thereby validating the excellent numerical accuracy of the spatio-temporal collocation method. Using the same approach, we also studied the stability of three-dimensional standing waves, and showed that they are unstable to certain two-dimensional harmonic disturbances whose frequencies are close to those of the base standing waves. Over long temporal periods, such an instability usually leads to quasi-periodic recurrences akin to those reported in [bryant1995water, zhu2003three].
In the future, we plan to exploit other problems regarding time-periodic water waves, including the stability and three-dimensional dynamics of standing waves, Faraday waves, and breathers. These problems require highly accurate numerical computations, hence a major challenge is to overcome the computer-memory limitation in our spatio-temporal collocation method. A promising approach is to replace the present Newton’s method by the Newton-Krylov method with preconditioners, which avoids explicitly constructing Jacobian matrix. Recently, this method has been successfully applied to the water-waves problems [pethiyagoda2014jacobian, ctugulan2022three]. Another challenge is to compute the time-periodic solutions of the full water-wave equations by combining the spatio-temporal collocation method and the boundary-integral method. This approach will allow computations in the strongly nonlinear regime where the quintic and other higher-order models would fail or become numerically inefficient. Recent advances on the Ewald summation technique in developing fast solvers of three-dimensional Laplace’s equation [ambrose2013small, duan2000ewald, fructus2005efficient] have shed light on this direction.

### v2

@@ 5 Temporal evolutions
$\widehat{p}=\widehat{\eta}+\frac{\mathrm{i}\widehat{\varphi}}{\sqrt{|\boldsymbol{k}|+1/|\boldsymbol{k}|}},\quad\widehat{q}=\widehat{\eta}-\frac{\mathrm{i}\widehat{\varphi}}{\sqrt{|\boldsymbol{k}|+1/|\boldsymbol{k}|}}.$ (57)
By combining (55) and (56), we have
@@
Figure 30: Temporal evolution of a non-resonant Case II standing wave $(k=l=1,H=0.29)$ perturbed by $0.001\cos(x+y)$. (a) Time histories of $|\widehat{p}_{1,1}|$ and $|\widehat{p}_{-1,1}|$. The black dashed lines label the instants when the curves reach maximums and minimums. (b-f) Representative surface profiles at $t=0,72T,137.1T,204.1T$ and $273.2T$.
@@
Figure 31: Temporal evolution of a non-resonant Case II standing wave $(k=l=1,H=0.5)$ perturbed by $0.001\cos(x+y)$. (a) Time histories of $|\widehat{p}_{1,1}|$ and $|\widehat{p}_{1,-1}|$. The black dashed lines label the instants when the curves intersect and reach maximums and minimums. (b-c) Surface profiles at $t=26T$ and $61.6T$.
@@
Next we examine the stability of three-dimensional standing waves. Zhu et al. [zhu2003three] have demonstrated that two-dimensional gravity standing waves, say, in $x$-direction with a wavenumber $\kappa$, are linearly unstable to harmonic disturbances $\mathrm{e}^{\mathrm{i}(\alpha x+\beta y)}$ if $\alpha^{2}+\beta^{2}\approx\kappa^{2}$. From a point of view of resonance, this condition implies that the disturbances have approximately the same frequency as the base standing waves, thereby supporting energy transfer between them. Over long periods, they showed that the instability leads to nearly cyclic return to the initial condition. Suppose a similar instability mechanism exists in gravity-capillary case, then the leading-order decomposition (1) implies that the three-dimensional standing waves are likely unstable to certain harmonic disturbances as well. We start to examine this idea by considering a non-resonant Case II standing wave with $k=l=1,H=0.29$, as shown in 30 (b). To satisfy the periodic boundary condition, we initially perturb this solution by a small disturbance $0.001\cos(x+y)$, which corresponds to two complex harmonics $0.0005\mathrm{e}^{\pm\mathrm{i}(x+y)}$. Figure 30 (a) shows the time histories of $|\widehat{p}_{1,1}|$ and $|\widehat{p}_{1,-1}|$ at integer temporal periods, which can be regarded as ensemble averages measuring the typical amplitudes of $\cos(x+y)$ and $\cos(x-y)$ modes. These two curves clearly demonstrates that energy is transferred between the two modes in a quasi-periodic fashion with an approximate period of $70T$. Therefore, the base standing wave is unstable to the initial perturbation. Panels (c-f) show four representative surface profiles at $t=71T,137.1T,204.1T$ and $273.2T$. Together with the initial profile in (b), these form a near cyclic return to the initial standing wave. For base standing waves with larger amplitude and the same initial disturbance, we observe similar cyclic returns occurring on relatively shorter temporal periods. Furthermore, there could be two alternatively appearing quasi-two-dimensional modes. A temporal evolution initiated with a Case II standing waves for $k=l=1,H=0.5$ is shown in figure 31. Panel (a) presents the curves of $|\widehat{p}_{1,1}|$ and $|\widehat{p}_{1,-1}|$ taken at each integer period, and panels (b) and (c) correspond to the surface profiles at $t=26T$ and $61.6T$ when the quasi-two-dimensional modes appear.
@@
Figure 32: Temporal evolution of a branch-$1$ standing wave $(k=1,l=1.5312,H=0.05)$ perturbed by $0.001\cos(x+1.5312y)$. (a) Time histories of $|\widehat{p}_{1,1.5312}|$, $|\widehat{p}_{1,-1.5312}|$, and $|\widehat{p}_{0,3.0624}|$. (b-c) Surface profiles at $t=0$ and $453.2T$.
@@
In general, consider three-dimensional standing waves dominated by $\cos(kx)\cos(ly)\cos(\omega t)+\cos(lx)\cos(ky)\cos(\omega t)$ and perturbed by $\delta\cos(kx\pm ly),\delta\ll 1$. It is found that such an instability is universal and the resulting energy-transfer rate is related to the frequencies of the base standing wave and the one used as initial perturbation. For two dimensional standing waves with wavenumber $\kappa=\sqrt{k^{2}+l^{2}}$, Concus [concus1962standing] showed that the frequency $\omega$ is a monotonically decreasing function of wave amplitude unless $\kappa\in(\kappa_{1},\kappa_{2})$. For infinite deep water, $\kappa_{1}=0.5154$ and $\kappa_{2}=\sqrt{2}/2$ (see figure 1 in [concus1962standing]). Therefore, if $\kappa$ lie outside this specific interval and the frequency curves of the base standing waves have down-shifts close to the bifurcation point, the base standing waves are likely unstable. Otherwise, there is no overlap between the the frequency curves of the base stadning waves and the perturbations, admitting no significant energy transfer. To confirm this, we consider the branch-$1$ standing waves with $k=1,l=1.5312$. The leading Fourier components can be decomposed into
@@
We first consider the initial disturbance $0.001\cos(x+1.5312y)$ whose oscillation frequency is close to $2.8188$. As shown in figure 22, the frequency curve of the base standing wave is slightly above this value for $H<0.12$. Figure 32 exhibits the temporal evolution of the standing wave with $H=0.05$ whose frequency is $\omega=2.8224$. Panel (a) shows the time histories of $|\widehat{p}_{1,1.5312}|,|\widehat{p}_{1,-1.5312}|$, and $|\widehat{p}_{0,3.0624}|$, which are taken at integer temporal periods and represent the typical amplitudes of $\cos(x+1.5312y)$, $\cos(x-1.5312y)$, and $\cos(3.0624y)$ modes. Energy transfer among the three modes is weak and only has influences over long periods. Note that the evolution also demonstrates a nearly periodic return to the initial state (with a phase shift) with an approximate period of $550T$, as shown by the surface profiles in panels (b) and (c).
@@ 6 Conclusions
We showed that the cubic and quintic models align closely with the full-potential formulation in both two and three dimensions. Typical non-resonant gravity-capillary standing waves exhibit round crests and troughs, in contrast to the pyramidal shape of pure gravity standing waves. When resonances occur, standing waves exhibit complicated and rapidly oscillating surface patterns, and their bifurcations could break into multiple, disjoint solution branches. In particular, we considered the resonant standing waves that support the three-wave resonance, generalising the concept of classical Wilton ripples occurring in two-dimensional gravity-capillary travelling waves. For both collinear and non-collinear resonant triads, we identified the existence of multiple solution branches bifurcating from the linear solutions, along with various resonant standing waves featuring square, hexagonal, and more complex flower-like surface patterns. We examined the temporal periodicity of the computed standing waves via initial-value calculations, validating the excellent numerical accuracy of the spatio-temporal collocation method. Using the same approach, we also studied the stability of three-dimensional standing waves, and showed that they could be unstable to certain harmonic disturbances whose frequencies are close to those of the base standing waves. Over long temporal periods, such an instability usually leads to the quasi-periodic return to the initial state with a possible phase shift, akin to those reported in [bryant1995water, zhu2003three].
In the future, we plan to exploit other problems regarding time-periodic water waves, including the stability and long-term dynamics of standing waves, Faraday waves, and breathers. These problems require highly accurate numerical computations, hence a major challenge is to overcome the computer-memory limitation in our spatio-temporal collocation method by using some matrix-free methods. Two promising approaches are the preconditioned Newton-Krylov method which has been successfully applied to several water-waves problems [pethiyagoda2014jacobian, ctugulan2022three], and the adjoint-based variational method which has been developed to find the time-periodic solutions of the Kuramoto–Sivashinsky equation [azimi2022constructing, ashtari2023jacobian], the Navier-Stokes equations [parker2022variational], the Benjamin-Ono equation [ambrose2010computationb], and the vortex-sheet problem [ambrose2010computationa]. Another challenge is to compute the time-periodic solutions of the full water-wave equations by combining the spatio-temporal collocation method and the three-dimensional boundary-integral method. This approach will allow computations in the strongly nonlinear regime where the quintic and other higher-order truncated models would fail or become numerically inefficient. Recent advances in developing fast solvers of three-dimensional Laplace’s equation using the Ewald summation technique [ambrose2013small, duan2000ewald, fructus2005efficient] have shed light on this direction.